MATLAB Assignment #1: Solve the following problems by writing commands in by writing a program in a script file and executing the file. Tum in a printout of the results and email the m file. All work should be individual, not shared
9. Let a 13, b= 4.2 (4b) abc d= ab+c then evaluate da - (a- )(c+ d) cb c+ d 10. A cube has a side of 18 cm. Detemine the radius of a sphere that has the same surface area as the cube. Determine the radius of a sphere that has the same volume as the cube a. b. 11. Let x 2 then evaluate 4 3r 4 x+1 12. Evaluate the trigonometric identity for 12° 4tanx-4tan'x 1-6tan2x+ tan'x 13. In the triangle shown a 5 in, b-7 in, and y- 25 a. Calculate the length of c by substituting the variables in the Law of Cosines. b. Calculate the ang les a and B (in degrees) using the Law of Sines
Answer #1

12)octave:2> x = (12*3.141592653589793/180) tan(4*x) ((4*(tan(x) )) - (4*tan(x) *tan(x)*tan(x) )) / (1- (6*tan(x)*tan(x))+ LHS =9)octave:2> a 13 = b 4.2 (4*b/a) (a b*c)/(a+b+c) ((a*b)/(c+d))+( (d*a)/(c*b))- ( (a- (b*b)) = = (c+d) e = 13 a b 4.2000 1.2923

0)a) Given (ube has ida o hence Side length a- 18 Cm Surface ahea of (ube =6a Sphere of hadius 4TTA Surkace Ceea Hence equeoctave:2> a = 18 r sqrt((6 a*a)/(4*3.14159265358979)) 18 a = 12.438 (OLcube hes side of 18cm 0)b) Gi ven hencede legth a 18 cm Volume ok cube =a Volume of Sphere ok ladiusy TI1 3 Hence equalion18 octave: 2> a r nthroot (((3*a*a*a)/(4*3.141592653589793)),3) 18 a 11.166 r